# Does quantum information require additional structure?

### v2

@@ Abstract
We consider interpretative problems of models of physical reality, including quantum information phenomenon in the context of Heisenberg’s classification of the fundamental physical theoretic models under the role of universal constants Planck’s constant $h$ and speed of light $c$. We introduce the correspondence principle between physical reality and mathematical models, and we discuss its significance in relation to physical models. We consider the status of quantum information in the standard quantum model, and based on the correspondence principle, we propose an interpretation of the wave function as a mathematical representation of quantum information. In this context, we consider Clauser’s analysis of incompatibility formulations of quantum theory in laboratory space and configuration space in the context of local realism. Then, we introduce the hypothesis of quantum space of directly unobserved relations, which precede correlations in the “classical” Minkowski space-time and are compatible with the Reichenbach common cause principle. Finally, we present the Chyliński model as an example of quantum relational space, which predicts potentially measurable effects for the bound states.
@@
“Heisenberg’s motivation for studying physics
was not only to solve particular problems but also to
illuminate the discussion of broad philosophical question”
@@ 1 Introduction
Quantum theory is the best theory we have to describe physical reality. “All our former experience with application of quantum theory seems to say: what is predicted by quantum formalism must occur in laboratory” [1]. It can be said that it works very well, which is why some physicists claim that “Quantum theory needs no interpretation” [2]. Nevertheless, there is a significant group of scientists for whom the lack of a coherent, commonly accepted interpretation is a source of discomfort. There have been many significant attempts to build a consistent interpretation, such as Copenhagen interpretation, Hidden Variable interpretations [3, 4, 5], and Relativistic interpretations [6, 7, 8]. However, there is still no consensus on what quantum mechanics is. There have been various propositions: “only on relative facts” [8], “only on correlations and only correlations” [9], “on the only fields” [10], “All is $\psi$[11], “The everything-is-a-quantum-wave” [12].
@@
One of the obstacles to a deeper understanding of quantum theory is that it radically deviates from classical patterns. Even pioneer quantum mechanics formulated nearly a hundred years ago, it appeared as a kind of quantum inscription encoded in the language of Hilbert space - a set of rules specifying how to make probabilistic predictions of results of future measurements in laboratories. The scale of difficulty in its deciphering reflects the fact that the discoveries of such phenomena as EPR correlations (entanglement), quantum computing [13, 14], quantum teleportation [15], existence the quantum correlations beyond entanglement [17, 16, 18] took place long after its birth. Quite surprasingly despite winning the battle with nature in an experimental demonstration of breaking Bell’s inequality (Nobel Prize in physics 2022) “…there it still no concensus about what such violations actually imply” [19] (and ref. there in). Moreover, quantum theory does not offer any simple footbridge from the quantum world to our actual world (see in this context [20, 21, 22, 23, 24, 25, 26, 27]. This issue covers many aspects: it has a long history and a vast literature, including the famous Wigner paradox [28, 29, 30].
Paradoxically, although the concept of quantum information lies at the heart of quantum theory, it has rather little impact on its interpretative problems. The vast majority of considerations about the meaning of quantum formalism do not take into account the problem of the relationship between physical reality and mathematical models, including the role of universal constants. This work does not aim to solve these drawbacks of quantum theory but rather attempt to consider the quantum inscription in the context proposed here the correspondence principle and the concept of quantum information.
@@
In sec.2, we introduce the correspondence principle between physical reality and mathematical models and we discuss its significance in relation to physical models (theories) in the context of their classification under the roles of universal constants. In sec.3, we briefly discuss the relativistic classical model and the standard (nonrelativistic) quantum model, emphasizing the roles of universal constants and symmetry groups. In sec.4, we consider the status of quantum information in the standard quantum model, and based on the correspondence principle, we propose an interpretation of the wave function as a mathematical representation of quantum information. In sec.5, we consider Clauser’s analysis of incompatibility formulations of quantum theory in laboratory space and configuration space in the context of local realism. In sec.6, we hypothesize that quantum information requires the existence of an intrinsic relational quantum space in which reversible premeasurements take place and which precedes the space of correlations actualized in Minkowski space. In sec.7, we present Chyliński’s model of quantum relational continuum space, which leads to measurable effects and, at the level of operational observation, is consistent with special relativity.
@@ 2 Heisenberg’s classification of fundamental models and the correspondence principle
Looking at the impressive development of physical sciences, a fundamental question constantly comes back to us: What is the reason why nature allows itself to be described by means of mathematical structures and why, out of many potential structures, it accepts only certain selected ones, e.g. Riemannian geometry but non-set-theoretic geometry with Banach-Tarski paradox [31]. This peculiar selectivity of nature can be naturally related to the existence of an active observer, equipped with mathematical tools, which organize his rational thinking about the world and allow him to create models of physical reality. By the term active observer, we mean the conscious observer endowed with free will who can have access to the system directly or indirectly via control and measurement interactions associated with an experimental setup. However, the action of an active observer is accompanied by the activity of nature itself due to the existence of fundamental interactions. Hence, the potential connection between the mathematical structure of models and the structure of the described reality is far from obvious [32, 33, 19]. Two approaches can be distinguished [33]: the ”theory-driven” approach, which assumes that the production of experimental data is governed by a pre-existing mathematical model and the ”data-driven” approach according to which experimental data are considered to be immutable facts, which require to construct corresponding mathematical models. In this context, it is tempting to introduce a correspondence principle (CP):
@@
Any consistent successful partial description of nature is a sort of correspondence between the structure of physical reality and the structure of its mathematical representation. Here, by the term “physical reality,” we mean a set of empirical observer-independent data, which includes all fundamental physical constants [34]. The term “partial description of nature” expresses the fact that, in practice, an observer has access to only a part of the world called a “system”. The term “correspondence” means that physical reality expressed in the language of specific mathematical structures is true according to the theory of reflection [35]. The problem is that the correspondence principle itself does not specify how elements of mathematical structure are related to physical reality. Therefore, the observer must somehow establish a set of interpretative statements to extract meaning from mathematical formalism. He can always, in spirit Gödel, supplement the model with additional structure before applying it [36].
@@
It is significant that “canonical” models of reality, including the Standard Model need constants parameters (fundamental constants) that cannot be explained within the framework of the given model, and they can only be determined from the experiment. Note that the correspondence principle allows for an “operational” interpretation of the model-independent parameters, as interaction constants between the mathematical structure of a model and the structure of the reality it describes. These constants impose global realistic constraints on the mathematical structure of models. There are the “conditional quantities” necessary for the model to reproduce the experimental results. In particular, they build relationships between quantities that were thought to be incommensurable.
@@
Note that there is no characterization of fundamental constants, and their list usually depends on what we accept as a fundamental model (theory). Hence, not all constants are on the same footing. Historically, the Planck’s constants $h$, and speed of light $c$ promoted to the role of universal constants. Here, we will use the term “universal constants” for which we assume that they meet the universality postulate: values of universal constants measured in any reference system always take the same values (see in this context[37, 38, 39]).
@@
According to Heisenberg, the introduction of universal constants led to abrupt changes in the paradigms of four fundamental models (theories) [40]: i) the non-relativistic classical (Newtonian model) (NRC): $01/c=$ 0, $h=0$, ii) relativistic classical (RC): $1/c\neq 0$ and $\hbar=0$, iii) the non-relativistic quantum (NRQ): $01/c=$ 0, $\hbar\neq 0$ iv) the relativistic quantum field (RQF): $1/c\neq 0$ and $\hbar\neq 0$. This classification does not take into account the gravitational constant $G$, which may be of fundamental importance in the quantum-to-classical transition [41].
@@
Note that there were also attempts to build a quantum relativistic models with $h/c\neq 0$ [42, 43, 44]. In the final part of the paper, we will present Chyliński’s relation-space-time model $h/c\neq 0$ [43] which, as far as we know, has not been considered so far, in any physical or philosophical context.
@@
3 Relativistic classical model
The RC based on the Galileo-Einstein principle and the observer-independent of the speed of light $c$ leads to the concept of Minkowskian space-time $M$ and its Poincare symmetry group as a relativistic model of space-time. Note that in the RC the measurement is passive in the sense that it does not disturb the system such that information can be freely copied by repeatedly interacting with the system. It can be explained by assuming that the observer only has at his disposal “classical” information carriers (CIC) which carry an arbitrarily small portion of energy and momentum. Hence, the recoil of the system being “measured” by the CIC is reduced to zero. It makes the RC non-contextual, as the reproducible properties of the system can be defined independently of any context. It should be noted here, however, that if one assumes that an active observer can freely influence physical systems without pre-defined correlations, this always generates contextuality [32].
@@ 3 Relativistic classical model
Note that the mathematical structure of the canonic phase space can be (according to CP principle) directly attributed to the physical structure as the probability distributions assigned to the classical phase space regions unambiguously correspond to the possible states of the system in which it has specific properties. Clearly, the RC is not compact. As Einstein already noted: “in the present state of theoretical physics, measures, and clocks must be considered as independent concepts, and we are far from a precise knowledge of theoretical foundations to be able to provide the structure of these objects.”
4 Non-relativistic (standard) quantum model
@@
In contrast to the classical models, the role of CP in relation to the quantum ones is much more subtle, as that NRQ strongly deviates from classical patterns. It appears rather as a quantum inscription written in the language of Hilbert space where the observables represented by Hermitian’s operators and the quantum states by positive unit trace matrices cannot be directly attributed to the physical reality. This step changes the paradigm in the description of nature has its origins in the fact that the mathematical structure of the NRQ includes Planck’s constant $h\neq 0$, which i) determines the ultimate limit on the non-commutativity of measurement operators that impose constraints on the possibility of extracting information by measurement from a quantum system[46, 47], and separation of the measured system (“specimen”) from the measuring device ceases to be obvious as in the RC. It implies ii) breaking the rescaling of the spatiotemporal extension of the systems by determining the finite volume of the phase cell, iii) the existence of the quantum information carriers (QIC),
The NRQ model applies to the phenomenological regime in which energies are low, and the effects of special relativity and high-energy physics can be neglected. The relativistic effects can be described within the framework of the RQ, which is based on the Lorentz-Poincare symmetry $M_{4}$. Note, however, that the NRQ model is fundamental because it is closed and compact. Closed theories in Galilean $G_{4}$ space-time can introduce hypothetical particles characterized by their absolute masses and spins [48]. In particular, the intrinsic symmetry of the non-relativistic Heisenberg’s, Schrödinger’s, and Pauli’s equations is also the symmetry of their empty Galilean group $G_{4}$ background. As one knows, the unitary representations of the $G_{4}$ have a nontrivial projective nature with physical consequences: Bargmann’s superselection mass rule [49], the noncommutativity of operators momentum and position, the possibility of a spin angular momentum (Pauli’s model). The “elementary” particles are defined by irreducible projective unitary representations of the Galilean group in some Hilbert space, which do not interact directly but via a field that mediates the interaction between particles [50, 49].
@@
It is generally accepted that the NRQ is neither universal nor the best model of reality, but it is certainly fundamental, and it can be a starting point for more advanced models such as RQ or S-matrix formalism. It is compact in the sense that it reproduces the extent of stable macroscopic measures with a hierarchical structure (quarks, nucleons, nuclei, atoms, molecules…). Since NRQ was completed, there have been times when physicists thought that its “quantum code” was deciphered completely, like the Rosetta Stone. Therefore, a big surprise was when Einstein Podolski and Rosen [51] discovered that NRQ predicts the existence of “spooky” correlations between properties of particles regardless of distance called by Schrödinger ”Verschrankung” (entanglement) [52]. Even though more than half a century has passed since its discovery, its full physical meaning still remains far from clear. Recently, it was argued that “…a violation of Bell-type inequalities by experimental data merely reflects the properties of the process, chosen by the experimenter, to select groups of data, not an intrinsic property of the data themselves.”[19]. Quantum entanglement and Born law break fundamental features of the RC: the separability and classical determinism, but the full significance of these facts seems to require new concepts.
Interestingly, the NRQ predicts a number of quantum phenomena without any reference to RQ: electron diffraction in the Davisson-Germer experiment, the quantum Hall effect, scattering experiments which involve collisions where radiative effects can be ignored [36]. However other well-known quantum phenomena, such as the two-slit experiment random decay times and neutron diffraction cannot be described completely within the framework of NRQ without referring to RQ if the detection processes are included.
5 Quantum model as a paradigm for quantum information
@@
There have been many obstacles to deciphering the potential physical implications of rich structure NRQ. It was only a series of groundbreaking theoretical and experimental papers showed that at a fundamental level, nature reveals the existence of a new, subtle property of physical systems - non-classical (or quantum) information [53], on the processing of which the mathematical structure of Hilbert space imposes strong constraints [54]. It cannot be cloned, broadcasted, or erased. It is nobody’s knowledge: it is neither epistemic nor ontological. It behaves as an independent physical entity. And what most important quantum information can be measured, processed (manipulated) stored [55, 56, 57, 1, 58, 59].
The discovery that quantum information is a new subtle resource for non-classical tasks [60, 61, 62, 63, 64, 65, 66, 67, 68, 69] revealed its operational meaning, which was of great importance for the development of quantum information theory and quantum technologies. One of the arguments in favor of the reality of quantum information is the resource criterion: If a quantum property associated with the system can be used as resource in non-classical tasks it is related to the physical reality [70].
However, the problem is that quantum information is relational in the sense it depends of the context [71, 72]. Hence, it has a “palette” of the properties within the different physical measurement contexts referred to methods of preparation (source information) and measurement contexts [73, 74].
On a formal level, composite quantum systems can be decomposed into subsystems in many different inequivalent ways. For example, if we decompose the Hilbert space $H$ of compound system S as a tensor product of smaller spaces $H=H_{1}\otimes H_{2},\ldots,\otimes H_{n}$ then the unit vectors and the states of the system can be very different, so different decompositions can have different properties. This difference is due to the fact that the symmetry group $U(d_{1})\times\ldots\times U(d_{k})$, $d_{1}=\dim(H_{i})$ is much smaller than $U(d)$, $d=\prod_{i=1}^{k}d_{i}$.The particular decomposition for the compound system represents meronomic reference frame. It may happen that in one meronomic system, separable (product) states may become entangled and vice versa. [75, 76]. It can be implemented in various possible ways, and an observer can have access to the system via control and measurement interactions associated with a corresponding experimental setup.
For quantized electromagnetic fields, the “quantum” and “classical” features of quantum information depend on the preparation of the initial states and the method of detection [77]. Namely, in contrast to classical communication where the information is encoded in classical properties of electromagnetic field (intensity or phase) quantum communication needs to send quantum information carriers such as nonclassical states of light e.g., squeezed states or Fock states that carry a well-defined number of photons. In the typical communications links, the task is an optimization over all ensembles of input quantum states used to carry information under relevant physical constraints. It is the Holevo theorem that provides a tight bound on the mutual information attainable for a given ensemble of input quantum states.
@@
Although the notion of quantum information was known before [78], it was only the concept of transmitting quantum states intact in the papers of the teleportation [15] and Schumacher quantum coding [79] that was revolutionary, because at that time the extremely epistemic Copenhagen’s interpretation was dominant. According to it, the wave function $\psi$ (quantum state) is merely the description of the measuring and preparing apparatuses. From such a point of view, it was rather impossible to think that the message could be something else than Shannon’s classical information. However, the correspondence principle prefers interpretation according to which: the wave function $\psi$ (the quantum state) is a mathematical representation (picture) of quantum information.
@@
Note that the above interpretation goes beyond the dilemma of the Scylla of ontology and the Charybdis of instrumentalism, and it is consistent with the well-known fact that in quantum models, the fundamental symmetry constraints are imposed on the quantum state rather than on probabilities.
To our knowledge, there no formal, commonly accepted definition of quantum information. Intuitively, one can think that the quantum information is what is carried by quantum systems in the given setup context. Nevertheless, there are fundamental “no go” broadcasting theorems [80, 81] that impose strong constraints on the processing of quantum information. More precisely, there are no completely positive trace-preserving maps that can duplicate any unknown quantum states. Perfect broadcasting is impossible because for the non-commuting states two copies (however correlated) will have more information than a single copy. For pure states, broadcasting is equivalent to cloning. The difference between cloning and broadcasting is that in cloning, one is required to obtain independent copies: In broadcasting, correlations between the copies are allowed [70].

### v3

@@ Abstract
We consider the status of quantum information in the quantum theory and based on the correspondence principle, we propose an interpretation of the wave function as a mathematical representation of quantum information. We consider Clauser’s analysis of incompatibility formulations of quantum theory in laboratory space and configuration-space in the context of local realism. Then, we introduce the hypothesis of quantum space of directly unobserved relations, which precede quantum correlations, and are compatible with the Reichenbach common cause principle. The possible implications of the hypothesis are discussed in the context of the latest experimental and theoretical results on the dynamics of entanglement formation in helium atoms. Finally, we present the Chyliński model as an example of quantum relational continuum space, which predicts potentially measurable effects for the bound states.
@@
“Heisenberg’s motivation for studying physics was not only to solve particularproblems but also to illuminate the discussion of broad philosophical question”
@@ 1 Introduction
Quantum theory is the best theory we have to describe physical reality. “All our former experience with the application of quantum theory seems to say: what is predicted by quantum formalism must occur in laboratory” [1]. It can be said that it works very well, which is why some physicists claim that “Quantum theory needs no interpretation” [2]. Nevertheless, there is a significant group of scientists for whom the lack of a coherent, commonly accepted interpretation is a source of discomfort. There have been many significant attempts to build a consistent interpretation, such as Copenhagen interpretation, Hidden Variable interpretations [3, 4], and Relativistic interpretations [5, 6, 7]. However, there is still no consensus on what quantum mechanics is. There have been various propositions: “only on relative facts” [7], “only on correlations and only correlations” [8], “on the only fields” [9], “All is $\psi$[10], “The everything-is-a-quantum-wave” [11]. One of the obstacles to a deeper understanding of quantum theory is that it radically deviates from classical patterns. Even pioneer quantum mechanics formulated nearly a hundred years ago, it appeared as a kind of quantum inscription encoded in the language of Hilbert space - a set of rules specifying how to make probabilistic predictions of results of future measurements in laboratories. The scale of difficulty in its deciphering reflects the fact that the discoveries of such phenomena as EPR correlations (entanglement), quantum computing [12, 13], quantum teleportation [14], existence of the quantum correlations beyond entanglement [16, 15, 17] took place long after its birth. Quite surprisingly despite winning the battle with nature in an experimental demonstration of breaking Bell’s inequality (Nobel Prize in physics 2022 [18]) “…there is still no consensus about what such violations actually imply” [19] (and ref. therein). Moreover, quantum theory does not offer any simple footbridge from the quantum world to our actual world (see in this context [20, 21, 22, 23, 24]. This issue covers many aspects: it has a long history and a vast literature, including the famous Wigner paradox [25, 26].
@@
Paradoxically, although the concept of quantum information lies at the heart of quantum theory, it has rather little impact on its interpretative problems. The vast majority of considerations about the meaning of quantum formalism do not take into account the problem of the relationship between physical reality and mathematical models, including the role of universal constants. This work does not aim to solve these drawbacks of quantum theory but rather attempts to consider the quantum inscription in light of the correspondence principle and the concept of quantum information. In Sec.2, we introduce the correspondence principle between physical reality and mathematical models and we discuss its significance in relation to physical models (theories) in the context of their classification under the roles of universal constants. In Sec.3, we briefly discuss the relativistic classical model and the standard (nonrelativistic) quantum model, emphasizing the roles of universal constants and symmetry groups. In Sec.4, we consider the status of quantum information in the standard quantum model, and based on the correspondence principle, we propose an interpretation of the wave function as a mathematical representation of quantum information. In Sec.5, we consider Clauser’s analysis of incompatibility formulations of quantum theory in laboratory space and configuration-space in the context of local realism. In Sec.6, we hypothesize that quantum information requires the existence of an intrinsic relational quantum space in which reversible pre-measurements take place and which precedes the space of correlations actualized in Minkowski space. In Sec.7, we present Chyliński’s model of quantum relational continuum space, which leads to measurable effects and, at the level of operational observation, is consistent with special relativity.
@@ 2 Heisenberg’s Classification of Fundamental Models and the Correspondence Principle
Looking at the impressive development of physical sciences, a fundamental question constantly comes back to us: What is the reason why nature allows itself to be described by means of mathematical structures and why, out of many potential structures, it accepts only certain selected ones, e.g., Riemannian geometry but non set‑theoretic geometry with Banach‑Tarski paradox [27]. This peculiar selectivity of nature can be naturally related to the existence of an active observer equipped with mathematical tools, which organize his/her rational thinking about the world and allow him/her to create models of physical reality. By the term active observer, we mean the conscious observer endowed with free will who can have access to the system directly or indirectly via control and measurement interactions associated with an experimental setup. However, the action of an active observer is accompanied by the activity of nature itself due to the existence of fundamental interactions. Hence, the potential connection between the mathematical structure of models and the structure of the described reality is far from obvious [28, 29, 19]. Two approaches can be distinguished [29]: the ”theory-driven” approach, which assumes that the production of experimental data is governed by a pre-existing mathematical model, and the ”data-driven” approach, according to which experimental data are considered to be immutable facts, which require to construct corresponding mathematical models. In this context, it is tempting to introduce a correspondence principle (CP):
{mdframed}
Any consistent successful partial description of nature is a sort of correspondence between the structure of physical reality and the structure of its mathematical representation.
@@
Here, by the term “physical reality”, we mean a set of empirical observer-independence data, which includes all fundamental physical constants [30]. The term “partial description of nature” expresses the fact that, in practice, an observer has access only to a part of the world called a “system”. The term “correspondence” means that physical reality expressed in the language of specific mathematical structures is true according to the Aristotle’s theory of truth [31, 32]. The problem is that the correspondence principle itself does not specify how elements of mathematical structure are related to physical reality. Therefore, the observer must somehow establish a set of interpretative statements to extract meaning from mathematical formalism. He can always, in the spirit of Gödel, supplement the model with additional structure before applying it.
@@
It is significant that “canonical” models of reality, including the Standard Model need constants parameters (fundamental constants) that cannot be explained within the framework of the given model, and they can only be determined from the experiment. Note that the correspondence principle allows for an “operational” interpretation of the model-independent parameters, as interaction constants between the mathematical structure of a model and the structure of the reality it describes. These constants impose global realistic constraints on the mathematical structure of models. These constants are necessary for the model to reproduce the experimental results. In particular, they build relationships between quantities that were thought to be incommensurable.
@@
Note that there is no characterization of fundamental constants, and their list usually depends on what we accept as a fundamental model (theory). Hence, not all constants are on the same footing. Historically, the Planck’s constants $h$, and speed of light $c$ were promoted to the role of universal constants. Here, we will use the term “universal constants” for which we assume that they meet the universality postulate: values of universal constants measured in any reference system always take the same values (see in this context[33, 34, 35]).
@@
According to Heisenberg, the introduction of universal constants led to abrupt changes in the paradigms of four fundamental models (theories) [36]: i) relativistic classical (RC): $1/c\neq 0$ and $\hbar=0$, ii) the non-relativistic quantum (NRQ): $01/c=$ 0, $\hbar\neq 0$ iii) the relativistic quantum field (RQF): $1/c\neq 0$ and $\hbar\neq 0$. This classification does not take into account the gravitational constant $G$, which may be of fundamental importance in the quantum-to-classical transition [37].
@@
Note that there were also attempts to build a quantum relativistic models with $\hbar/c\neq 0$ [38, 39, 40]. In the final part of the paper, we will present Chyliński’s relation-space-time model [39] which, as far as we know, has not been considered so far, in any physical or philosophical context.
@@
3 Relativistic Classical Model Versus Quantum Nonrelativistic model
Obviously, every one of the above models is a kind of idealization, while its predictive power is determined by a specific paradigm intimately related to universal constants. Specifically, the RC model ($\frac{1}{c}\neq 0$, $\hbar=0$) based on the Galileo-Einstein principle and the postulate that the speed of light in vacuum is the same for all observers leads to the concept of Minkowskian space-time $M$ and its Poincare symmetry group as a classical relativistic model of space-time.
@@ 3 Relativistic Classical Model Versus Quantum Nonrelativistic model
Note that according to CP, the mathematical structure of the canonic phase space can be directly attributed to the physical structure as the joint probability distributions assigned to the classical phase space regions unambiguously correspond to the possible states of the system in which it has specific properties. Thus, by knowing the state of the system we can say that we know the physical reality regardless of the measurement context. The later is only auxiliary tool for measurements that do not disturb the system. Thus information can be freely copied by repeatedly interacting with the system. Clearly, the RC is not compact. As Einstein already noted: “in the present state of theoretical physics, measures, and clocks must be considered as independent concepts, and we are far from a precise knowledge of theoretical foundations to be able to provide the structure of these objects.”
In contrast to the RC model, the role of CP in relation to the NRQ model is much more subtle, as the latter strongly deviates from classical patterns. It appears rather as a quantum inscription written in the language of Hilbert space where the observables represented by Hermitian’s operators and the quantum states by positive unit trace matrices cannot be directly attributed to the physical reality. Here, the measurement is active: It can change state and can also be part of quantum operation. Therefore, in contrast to the RC model the measurement context cannot be ignored. What most differs the NRQ model from RC one is quantum principle of superposition, which has quantum origin.
@@
This step change of the paradigm in the description of nature has its origins in the fact that the mathematical structure of the NRQ includes Planck’s constant $\hbar\neq 0$ implying i) the ultimate limit on the non-commutativity of measurement operators that impose constraints on the possibility of extracting information by measurement from a quantum system[42, 43], so separation of the measured system from the measuring device ceases to be obvious as in the RC, ii) breaking the rescaling of the spatiotemporal extension of the systems by determining the finite volume of the phase cell, iii) the existence of the quantum information carriers (QIC). The NRQ model has an amazing property; even though it does not include the speed of light $c$, it is consistent with the results of the RC model such as the impossibility of superluminal communication.
@@
The NRQ model applies to the phenomenological regime in which energies are low, and the effects of special relativity and high-energy physics can be neglected. The relativistic effects can be described within the framework of the RQ, which is based on the Lorentz-Poincare symmetry $M_{4}$. Note, however, that the NRQ model is fundamental because it is closed and compact. Closed theories in Galilean $G_{4}$ space-time can introduce hypothetical particles characterized by their absolute masses and spins. In particular, the intrinsic symmetry of the non-relativistic Heisenberg’s, Schrödinger’s, and Pauli’s equations is also the symmetry of their empty Galilean group $G_{4}$ background. As one knows, the unitary representations of the $G_{4}$ have a non-trivial projective nature with physical consequences: Bargmann’s superselection mass rule [44], the noncommutativity of operators momentum and position, the possibility of a spin angular momentum (Pauli’s model). The “elementary” particles are defined by irreducible projective unitary representations of the Galilean group in some Hilbert space, which do not interact directly but via a field that mediates the interaction between particles [44, 45]. It is generally accepted that the NRQ is neither universal nor the best model of reality, but it is certainly fundamental, and it can be a starting point for more advanced models such as RQ or S-matrix formalism. It is compact in the sense that it reproduces the extent of stable macroscopic measures with a hierarchical structure (quarks, nucleons, nuclei, atoms, molecules…). Since NRQ was completed, there have been times when physicists thought that its “quantum code” was deciphered completely, like the Rosetta Stone. Therefore, it was a big surprise when Einstein Podolski and Rosen[46] discovered that NRQ predicts the existence of “spooky” correlations between properties of particles regardless of distance called entanglement [47]. Even though more than half a century has passed since its discovery, its full physical meaning still remains far from clear. Recently, it was argued that “… a violation of Bell-type inequalities by experimental data merely reflects the properties of the process, chosen by the experimenter, to select groups of data, not an intrinsic property of the data themselves.”[19]. Quantum entanglement and Born law break fundamental features of the RC: the separability and classical determinism, but the full significance of these facts seems to require new concepts.
4 Quantum Information Puzzle
@@
There have been many obstacles to deciphering the potential physical implications of rich structure NRQ. It was only a series of groundbreaking theoretical and experimental papers showed that at a fundamental level, nature reveals the existence of a new, subtle property of physical systems - quantum information with no classical analog [48]. Quantum entanglement, the most non-classical feature of quantum information, manifests itself at both macro [49] and the micro scale [50, 51, 52]. Entanglement can be measured, processed, manipulated, and stored [53, 54, 49, 1, 55, 56]. The discovery that quantum information is a new subtle resource for non-classical tasks [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] revealed its operational meaning, which was of great importance for the development of quantum information theory and quantum technologies.
@@
Although the notion of quantum information was known before [67], it was only the concept of intact transmitting quantum states in the papers of the teleportation [14] and Schumacher quantum coding [68] that was revolutionary because, at that time, the extremely epistemic Copenhagen’s interpretation was dominant. According to it, the wave function $\psi$ (quantum state) is merely the description of the measuring and preparing apparatuses. From such a point of view, it was rather impossible to think that the message could be something else than Shannon’s classical information. Since $\psi$ is a central concept in the structure of quantum formalism the correspondence principle allows the interpretation of the $\psi$ wave function as a mathematical representation (image) of the quantum information.
@@
Importantly, the above interpretation (called here correspondence interpretation) goes beyond the dilemma of the Scylla of ontology and the Charybdis of instrumentalism, and it is consistent with the well-known fact that in quantum models, the fundamental symmetry constraints are imposed on the quantum state rather than on probabilities.
To our knowledge, there no formal definition of quantum information. Intuitively, one can think that the quantum information is what is carried by quantum systems. The idea that a concept of quantum information should be regarded as a fundamental ingredient in quantum physics has been proposed [48, 69] from a different points of view. In [69] the argument was mainly based on the concept of quantum information isomorphism, according to which the quantum description of nature is isomorphic to nature itself. Here we extend this approach by the correspondence principle (CP), which goes beyond a purely mathematical context.